# Optimal Bound for PCA with Outliers using Higher-Degree Voronoi Diagrams

### v2

@@ 2.2 Voronoi diagrams
Figure 2: Voronoi diagram of the furthest subspace of the 2nd-degree for 8 points in the two-dimensional space on one-dimensional subspaces. The projection of subspaces for the Voronoi diagram of the furthest subspace is shown on the unit circle (smaller circle) and the image of subspaces for the Voronoi diagram of the furthest subspace for the second furthest subspace is marked on the circle with radius 2 (larger circle). Space coloring indicates different cells and subspaces of the same color indicate the same cell. For example, in the above figure, 7 cells are specified, considering that 4 of these eight points belong to exactly one subspace, the number of possible cells is far less than all the possible combinations as outlier data.
@@ 3 Optimal PCA with Outliers
In this section, we present our algorithm for solving the PCA with Outliers problem in $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$ time. The solution involves obtaining a better subspace embedding for a matrix by removing its outliers. The challenge is that we do not know the actual inlier matrix to compute the scores, as an arbitrary set of rows of the given matrix might be outliers. Nevertheless, we can narrow down the available choices such that no possibility for a given combination of outliers, exists.
@@
Note that by proposition 3, if such an algorithm exists, it should be optimal. In each step, we pick a possible $\mathtt{k}$-tuple of data and label them as outliers, then compute PCA for the rest and measure the relevant error, and finally, depending on the least measured error, we get the output. Nevertheless, the actual problem is the number of possible $\mathtt{k}$-tuples of outliers. In order to specify each of these possible $\mathtt{k}$-tuples, we need to find their corresponding Voronoi cell. However, there is no need to obtain the exact Voronoi diagram as we only need to know if a set of data can be grouped as outliers, which can be determined by the emptiness of the corresponding cell. Algorithm 1 illustrates the above steps.
@@
1:procedure PCA($X_{n\times d}$,r,k)
2:     Compute the $\mathtt{(n-k)}$th-degree Voronoi diagram of the columns of $X_{n\times d}$ according to def 5
@@
4:     $\hat{\textbf{L}}\leftarrow$ corresponding $k$-tuples of cells in L
5:     $\text{solution}\leftarrow\arg\max_{S\in\hat{\textbf{L}}}\text{PCA}(X\setminus S,r,k)$
6:     return solution
7:end procedure
@@ 4 Randomized Algorithm
In this section, we present a randomized algorithm for solving the PCA with Outliers problem. We begin by introducing the concept of the $\alpha$-gap constant, which is a crucial component of our algorithm.
@@
($\alpha$-Gap Constant) For a set $\text{X}\subset\mathbb{R}^{d}$, there exists an $\alpha>0$ such that for an optimal solution of PCA with Outliers, involving a reduced matrix $L$ and a set of outliers $S$, we have for all $s\in S$ and $x\in\text{X}\setminus S$,
@@
Let $L$ be the optimal subspace for this problem. Since this is the optimal subspace for the inliers, any point classified as an outlier must have a greater distance to $L$ compared to the inliers. Otherwise, if an outlier was closer to $L$ than some inliers, it would be considered part of the main structure captured by $L$ rather than an outlier. Thus define $d_{1}$ as the distance of the $(n-k)$th farthest vector to $L$ and $d_{2}$ as the distance of the $(n-k+1)$th farthest vector to $L$ (i.e., the farthest inlier and the nearest outlier). We can define $\alpha$ as the ratio between these two distances:
@@
Using this constant, we provide a lower bound on the probability that a Voronoi cell contains the optimal subspace. This demonstrates that sampling multiple subspaces uniformly can yield the optimal result with high probability.
Next, we provide background information on Grassmannian manifolds and subspace projections to lay the structure of our algorithm. We prove a lemma concerning the distance between subspaces in a real vector space and how to extract an upper bound on the distance of one point to these subspaces. This allows us to apply a concept similar to the triangle inequality to the set of subspaces and vectors.
@@
Let $U$ and $V$ be two subspaces of the same rank in an $d$-dimensional real vector space, and let $x$ be a vector in this space. If the distance between $U$ and $V$ is $\mathtt{dist}(U,V)$(according to the Grassmannian metric), and the distance between $x$ and $U$ is $\mathtt{dist}(x,U)$, then the maximum distance of $x$ from $V$ is $\mathtt{dist}(x,U)(1+\mathtt{dist}(U,V))$.
@@
Let $P_{U}(x)$ and $P_{V}(x)$ denote the orthogonal projections of $x$ onto the subspaces $U$ and $V$, respectively. By definition, the distance from $x$ to $U$ is given by:
$\mathtt{dist}(x,U)=\|x-P_{U}(x)\|.$
The Grassmannian distance between the subspaces $U$ and $V$ ($\mathtt{dist}(U,V)$) is related to the largest principal angle $\theta_{\max}$ between them, such that:
$\mathtt{dist}(U,V)=\sin\theta_{\max}.$
To find the maximum distance from $x$ to $V$, consider the point $u=P_{U}(x)$. The distance from $u$ to $V$ is maximized when we project $u$ onto $V$ in the direction corresponding to $\theta_{\max}$. Thus, the distance from $u$ to its projection onto $V$ can be expressed as:
$\|u-P_{V}(u)\|=\|u\|\sin\theta_{\max}=\mathtt{dist}(x,U)\sin\theta_{\max}=\mathtt{dist}(x,U)\mathtt{dist}(U,V).$
By the triangle inequality, the distance from $x$ to $V$ can then be bounded as:
$\|x-P_{V}(x)\|\leq\|x-P_{U}(x)\|+\|P_{U}(x)-P_{V}(u)\|.$
Substituting the known distances, we get:
$\|x-P_{V}(x)\|\leq\mathtt{dist}(x,U)+\mathtt{dist}(x,U)\mathtt{dist}(U,V)=\mathtt{dist}(x,U)(1+\mathtt{dist}(U,V)).$
We now give another lemma regarding the maximal ball around a subspace, which preserves the relative distances of vectors from that subspace in the Grassmannian manifold.
@@
Let $x$ and $y$ be two vectors in an $d$-dimensional real vector space, and let $V$ be a rank $r$ subspace of this space such that $(1+\alpha)\cdot\mathtt{dist}(x,V)\leq\mathtt{dist}(y,V)$ for some $\alpha>0$. Then, the maximal open ball in the Grassmannian manifold $\text{Gr}(r,d)$ centered around $V$ such that for every subspace $U$ in this ball, $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$ has radius $\alpha$.
@@
Consider the Grassmannian manifold $\text{Gr}(r,d)$, which is the set of all $r$-dimensional subspaces of $\mathbb{R}^{d}$. Let $\mathtt{dist}(x,U)$ denote the Euclidean distance from a vector $x$ to a subspace $U$. Given the subspace $V$ and vectors $x$ and $y$ such that $(1+\alpha)\cdot\mathtt{dist}(x,V)\leq\mathtt{dist}(y,V)$, we aim to find the maximal open ball around $V$ within which for every subspace $U$, the inequality $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$ holds.
The distance between subspaces in the Grassmannian manifold is related to the principal angles between them. For any subspace $U$ close to $V$, the distances $\mathtt{dist}(x,U)$ and $\mathtt{dist}(y,U)$ can be bounded by Lemma 3. Specifically, for a subspace $U$ within a distance $\mathtt{dist}(U,V)$ in the Grassmannian from $V$, the distances are given by:
$\mathtt{dist}(x,U)\leq\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))$
and
$\mathtt{dist}(y,U)\leq\mathtt{dist}(y,V)(1+\mathtt{dist}(U,V)).$
To ensure $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$, we require:
$\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))\leq\mathtt{dist}(y,V)(1+\mathtt{dist}(U,V)).$
Given the condition $\mathtt{dist}(y,V)\geq(1+\alpha)\mathtt{dist}(x,V)$, substituting this into the inequality, we get:
$\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))\leq(1+\alpha)\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V)).$
Dividing through by $\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))$ (which is positive since $\mathtt{dist}(U,V)\geq 0$ and $\mathtt{dist}(x,V)\geq 0$), we obtain:
$1\leq 1+\alpha.$
Therefore, for the inequality to hold for all subspaces $U$ within this distance, the maximum $d$ must satisfy:
$1+\mathtt{dist}(U,V)\leq 1+\alpha.$
Hence,
$\mathtt{dist}(U,V)\leq\alpha.$
Thus, the maximal open ball in the Grassmannian manifold $\text{Gr}(r,d)$ centered around $V$ such that for every subspace $U$ in this ball, $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$ has radius $\alpha$. ∎
Now we explain how to compute the measure of the aforementioned maximal ball in the Grassmannian manifold $\text{Gr}(r,d)$. Since the Grassmannian manifold is endowed with the Haar measure, the measure of the maximal ball is simply the volume of the ball. Given that the radius of the maximal ball is $\alpha$, and the dimension of the Grassmannian manifold is $r(d-r)$, the measure of the ball can be computed using the formula for the volume of a ball in a Riemannian manifold.
In $\text{Gr}(r,d)$, the volume of a ball of radius $\alpha$ centered at a point $V$ is given by:
$\text{Vol}(\text{Ball}(V,\alpha))=C\times\alpha^{r(d-r)},$
where $C$ is a constant depending on the geometry of the Grassmannian manifold. As computing the exact constant $C$ for the volume of a ball in the Grassmannian manifold $\text{Gr}(r,d)$ is quite complex, we provide the next lemma to give a lower bound on this value.
@@
Let $\text{Gr}(r,d)$ be the Grassmannian manifold of $r$-dimensional subspaces of an $d$-dimensional real vector space. Consider a maximal open ball in $\text{Gr}(r,d)$ centered at a subspace $V$ with radius $\alpha$. The volume of such a ball can be expressed in terms of the volume of the entire Grassmannian and a proportionality constant that accounts for the curvature and dimensionality of the space.
Let $\text{Gr}(r,d)$ be the Grassmannian manifold of $r$-dimensional subspaces of a $d$-dimensional real vector space. Consider a maximal open ball in $\text{Gr}(r,d)$ centered at a subspace $V$ with radius $\alpha$. The ratio of the volume of such a ball to the volume of the entire Grassmannian manifold satisfies the following lower bound:
$\frac{\text{Vol}\left(\text{Ball}(V,\alpha)\right)}{\text{Vol}\left(\text{Gr}(r,d)\right)}\geq\alpha^{r(d-r)}\left(\frac{2e}{r(d-r)}\right)^{r(d-r)/2}.$
@@
The Grassmannian manifold $\text{Gr}(r,d)$ represents all $r$-dimensional subspaces of $\mathbb{R}^{d}$ and has a well-defined volume due to its compactness. The volume of the entire Grassmannian, denoted as $\text{Vol}(\text{Gr}(r,d))$, is given by:
$\text{Vol}(\text{Gr}(r,d))=\frac{\prod_{i=1}^{r}\left(\pi^{(d-i)/2}/\Gamma((d-i)/2+1)\right)}{\prod_{i=1}^{r}\left(\pi^{(r-i)/2}/\Gamma((r-i)/2+1)\right)\prod_{i=1}^{d-r}\left(\pi^{(d-r-i)/2}/\Gamma((d-r-i)/2+1)\right)}.$
To approximate the volume of a ball of radius $\alpha$ in $\text{Gr}(r,d)$ centered at a subspace $V$, we can consider the tangent space at $V$. For small $\alpha$, the ball in $\text{Gr}(r,d)$ behaves similarly to a Euclidean ball in this tangent space, whose dimension is $r(d-r)$. Thus, the volume of such a ball can be expressed as:
$\text{Vol}(\text{Ball}(V,\alpha))=C\alpha^{r(d-r)},$
where $C$ is a constant that depends on the geometry of the Grassmannian manifold.
To estimate the constant $C$, we consider the relationship between the volumes of the orthogonal groups and the structure of the Grassmannian. A lower bound for $C$ can be derived using the known volumes of these spaces. Specifically, the volume of a Euclidean ball in a $r(d-r)$-dimensional space is given by:
$\text{Vol}(\text{Ball}(V,\alpha))=\frac{\left(2\pi\right)^{r(d-r)/2}}{\Gamma\left(\frac{r(d-r)}{2}+1\right)}\alpha^{r(d-r)}.$
Therefore, the measure of the maximal ball of radius $\alpha$ in the Grassmannian manifold $\text{Gr}(r,d)$ can be bounded below by:
$\text{Vol}(\text{Ball}(V,\alpha))\geq\frac{\left(2\pi\right)^{r(d-r)/2}}{\Gamma\left(\frac{r(d-r)}{2}+1\right)}\alpha^{r(d-r)}.$
To further refine this bound, we apply Stirling’s approximation for the Gamma function:
$\Gamma(z)\approx\sqrt{2\pi}\left(\frac{z}{e}\right)^{z}.$
Using this approximation, we get:
$\Gamma\left(\frac{r(d-r)}{2}+1\right)\approx\sqrt{2\pi}\left(\frac{r(d-r)}{2e}\right)^{r(d-r)/2+1}.$
Substituting this into our volume expression, we obtain:
$\text{Vol}(\text{Ball}(V,\alpha))\geq\frac{\left(2\pi\right)^{r(d-r)/2}}{\sqrt{2\pi}\left(\frac{r(d-r)}{2e}\right)^{r(d-r)/2+1}}\alpha^{r(d-r)}=\frac{2^{r(d-r)/2}\alpha^{r(d-r)}}{\left(\frac{r(d-r)}{2e}\right)^{r(d-r)/2+1}}.$
Simplifying this expression:
$\text{Vol}(\text{Ball}(V,\alpha))\geq\frac{2^{r(d-r)/2}\alpha^{r(d-r)}(2e)^{r(d-r)/2+1}}{(r(d-r))^{r(d-r)/2+1}}.$
Thus, the simplified ratio of the measure of the maximal ball to the volume of the entire Grassmannian is:
$\mathbb{P}\left(\text{Ball}(V,\alpha)\right)\geq\alpha^{r(d-r)}\left(\frac{2e}{r(d-r)}\right)^{r(d-r)/2}.$
This provides a clear and robust lower bound for the measure of a maximal ball in the Grassmannian manifold $\text{Gr}(r,d)$. ∎
Now we give the following theorem which computes the measure of the Voronoi cell of an optimal subspace in the PCA with the Outliers problem, to be used later in our algorithm analysis.
@@
Let $X\subset\mathbb{R}^{d}$ be a set of vectors in the $d$-dimensional real vector space, and let $L$ be the optimal subspace of rank $k$ for the PCA with Outliers problem. If $VOD(L)$ is the Voronoi cell of this given subspace according to Definition 5, then selecting a rank $k$ subspace uniformly at random has a positive probability of at least:
$\mathbb{P}\left(VOD(L)\right)\geq\alpha^{r(d-r)}\left(\frac{2e}{r(d-r)}\right)^{r(d-r)/2}.$

### v3

@@ 3 Voronoi diagrams
Figure 2: 2nd-degree Voronoi diagram of the furthest subspace for 8 points in $\mathbb{R}^{2}$ space which can be used to solve PCA with outliers with goal dimension 1 and 2 outliers. The projection of subspaces for each Voronoi cell of the furthest subspace is shown on the unit circle (smaller circle) and the image of subspaces for the Voronoi cell of the second furthest subspace is shown on the circle with radius 2 (larger circle). Space coloring indicates different 2nd-degree cells and subspaces of the same color indicate the same cell. For example, in the above figure, 7 cells are specified, considering that 6 of these eight points belong to exactly one subspace, the number of possible cells is far less than all the possible combinations as outlier data.
Also, note that we use the variable $k$ as the number of outliers and the order and degree of Voronoi diagrams interchangeably since the latter is going to denote the number of given outliers.
@@ 4 Optimal PCA with Outliers
In this section, we present our exact algorithm for solving the PCA with Outliers problem in $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$ time. The solution involves obtaining a better subspace embedding for a matrix by removing its outliers. The challenge is that we do not know the actual inlier matrix to compute the scores, as an arbitrary set of rows of the given matrix might be outliers. Nevertheless, we can narrow down the available choices such that no possibility for a given combination of outliers, exists.
@@
Note that by proposition 3, if such an algorithm exists, it should be optimal. In each step, we pick a possible $\mathtt{k}$-tuple of data and label them as outliers, then compute PCA for the rest and measure the relevant error, and finally, depending on the least measured error, we get the output. Nevertheless, the actual problem is the number of possible $\mathtt{k}$-tuples of outliers. In order to specify each of these possible $\mathtt{k}$-tuples, we need to find their corresponding Voronoi cell. However, there is no need to obtain the exact Voronoi diagram as we only need to know if a set of data can be grouped as outliers, which can be determined by the emptiness of the corresponding cell. We choose non-empty cells because when a cell with index tuple $S$ is empty, there is no corresponding subspace with rank $r$ and $S$ as it $\mathtt{(n-k)}$-ordered nearest neighbors to form the required set. Algorithm 1 illustrates the above steps.
@@
1:$X$ ($n$ vectors in $\mathbb{R}^{d}$), $k$ (# of outliers), $r$ (Goal dimension)
2:Compute the $\mathtt{(n-k)}$-degree Voronoi diagram of the rows of $X_{n\times d}$ using theorem 1.
@@
[…]
7:for $\hat{S}\in\textbf{Index}$ do
8:  if $\text{loss}(\text{PCA}(X\setminus\hat{S},r),X\setminus\hat{S})\leq l$ then
[…]
10:   $S\leftarrow\hat{S}$
11:  end if
12:end for
13:return $\text{PCA}(X\setminus S,r)$
@@ 5 Randomized Algorithm
In this section, we demonstrate that, rather than explicitly constructing and enumerating every Voronoi cell—a process that can be prohibitively complex—one may instead perform random sampling over candidate subspaces and identify the optimal cell directly. By leveraging the intrinsic concentration phenomena of high-dimensional geometry, we establish sufficient conditions under which this randomized approach locates the correct Voronoi region with similar time complexity, resolving the need for full diagrammatic enumeration.
We begin by restating the setting of the problem. Let $X\subset\mathbb{R}^{d}$ be a set of $n$ points contained in the unit ball, and let $S^{*}\subset X$ be a distinguished subset of $k$ points (The set of optimal outliers). For any vector $v\in\mathbb{R}^{d}$ and an $r$-dimensional subspace $L\subset\mathbb{R}^{d}$, define
$\mathtt{dist}(v,L)=\|v-P_{L}v\|,$
where $P_{L}$ denotes the orthogonal projection onto $L$. Suppose that $L$ minimizes
$\sum_{p\in X\setminus S^{*}}\mathtt{dist}(p,L).$
To tackle this problem efficiently, we assume that the data points in $X\setminus S^{*}$ and the outliers in $S^{*}$ are well-separated. Consequently, we impose a separation condition, formalized in Lemma 1 to capture the property that outlier points are further from the best-fit subspace than inlier points. Which asserts the existence of a constant $\alpha>0$ such that for every $p\in X\setminus S^{*}$ and $q\in S^{*}$, the distance to the subspace $L$ satisfies the inequality: $\mathtt{dist}(q,L)\geq(1+\alpha)\cdot\mathtt{dist}(p,L).$ This condition ensures that the optimal subspace $L$ is not significantly influenced by the outliers, making the problem well-posed and enabling us to robustly computes the principal subspace. The separation condition is crucial for both the theoretical guarantees of our algorithm and the subsequent analysis.
@@
($\alpha$-Gap Constant) For a set $\text{X}\subset\mathbb{R}^{d}$, there exists an $\alpha>0$ such that for an optimal solution of PCA with Outliers, involving a reduced subspace $L$ and a set of outliers $S$, we have for all $s\in S$ and $x\in\text{X}\setminus S$,
@@
Let $L$ be the optimal subspace for this problem. Since this is the optimal subspace for the inliers, any point classified as an outlier must have a greater distance to $L$ compared to the inliers. Otherwise, if an outlier was closer to $L$ than some inliers, it would be considered part of the main structure captured by $L$ rather than an outlier. Thus define $d_{1}$ as the distance of the (nk)(n-k)th farthest vector to $L$ and $d_{2}$ as the distance of the (nk+1)(n-k+1) farthest vector to $L$ (i.e., the farthest inlier and the nearest outlier). We can define $\alpha$ as the ratio between these two distances:
@@
The $\alpha$-gap constant is fundamental in robust Principal Component Analysis (PCA) with outliers, as it guarantees a quantifiable separation between inliers and outliers. This separation is crucial for distinguishing genuine data points from anomalies and thus enhances the robustness of PCA algorithms. In particular, the $\alpha$-gap constant generalizes the ideal setting of Robust Subspace Recovery, where inliers lie exactly on a low-dimensional subspace [23], by allowing inliers to reside near the subspace under bounded noise while outliers remain sufficiently distant. Similarly, in probabilistic models such as Robust Probabilistic PCA [3], inliers drawn from a light-tailed distribution cluster closely around the subspace, whereas outlier distances are amplified in high dimensions.
From a computational perspective, the $\alpha$-gap constant facilitates the design of fixed-parameter tractable algorithms for PCA with outliers [13]. This theoretical soundness translates into practical relevance in applications such as medical imaging and fraud detection, where empirical evidence often suggests that outliers are distinctly separated from inliers (typically with $\alpha\geq 0.5$) [8].
Also, in high-dimensional spaces, the concentration of measure phenomenon forces inlier distances to the optimal subspace to tightly cluster around a mean value, while outlier distances tend to scale with the ambient dimension. Thus, although the curse of dimensionality is often viewed as a challenge, it in fact reinforces the $\alpha$-gap assumption by amplifying the separation between inliers and outliers [30, 13]. We are going to leverage this fact to establish lemma 4.
The following lemma establishes a lower bound on the measure of the set of subspaces $\hat{L}$ that are close to a fixed subspace $L$. This measure, quantifies the volume of subspaces $\hat{L}$ satisfying $\|P_{\hat{L}}-P_{L}\|\leq\epsilon$. This result is crucial because it allows us to reason about the probability that a randomly chosen subspace $\hat{L}$ will lie within a small neighborhood of the optimal subspace $L$.
@@
For a fixed subspace $L$ of rank $r$ and $\epsilon>0$ sufficiently small, define
$B_{\epsilon}(L)=\{\text{subspace $\hat{L}$ of rank $r$}:\|P_{\hat{L}}-P_{L}\|\leq\epsilon\}.$
Then there exists a constant
$c_{0}(d,r)=J_{0}\,\frac{\pi^{r(d-r)/2}}{\Gamma\Bigl(\frac{r(d-r)}{2}+1\Bigr)}\,c_{2}^{-r(d-r)},$
depending only on $d$ and $r$, such that
$\mu\bigl(B_{\epsilon}(L)\bigr)\geq c_{0}(d,r)\,\epsilon^{r(d-r)}.$
@@
A standard local coordinate chart for Grassmanian manifold near $L$ is given by representing any nearby subspace $\hat{L}$ as the graph of a linear map $A:L\to L^{\perp},$ where $L^{\perp}$ is the (dr)(d-r)-dimensional orthogonal complement of $L$. This identifies a neighborhood of $L$ with an open subset of $\mathbb{R}^{N}$, where N=r(dr).N=r(d-r). By perturbation results, the Davis–Kahan $\sin\Theta$” theorem [14] and its extension by Stewart–Sun [29], one obtains that for any linear map $A:L\to L^{\perp}$ with operator norm $\|A\|$ sufficiently small, there are constants $c_{1},c_{2}>0$, depending only on $d$ and $r$, such that
$c_{1}\,\|A\|\;\leq\;\bigl\|P_{\hat{L}}-P_{L}\bigr\|\;\leq\;c_{2}\,\|A\|,$
Thus, $\|P_{\hat{L}}-P_{L}\|\leq\epsilon$ implies $\|A\|\leq\epsilon/c_{2}$. The Euclidean volume in $\mathbb{R}^{N}$ of the ball of radius $\epsilon/c_{2}$ is
$\operatorname{Vol}\Bigl(B_{\epsilon/c_{2}}(0)\Bigr)=\frac{\pi^{N/2}}{\Gamma\Bigl(\frac{N}{2}+1\Bigr)}\Bigl(\frac{\epsilon}{c_{2}}\Bigr)^{N}.$
If the Jacobian of the coordinate chart is bounded below by $J_{0}>0$, then the measure of $B_{\epsilon}(L)$ satisfies
$\mu\bigl(B_{\epsilon}(L)\bigr)\geq J_{0}\,\frac{\pi^{r(d-r)/2}}{\Gamma\Bigl(\frac{r(d-r)}{2}+1\Bigr)}\,c_{2}^{-r(d-r)}\,\epsilon^{r(d-r)}.$
Thus, by setting
$c_{0}(d,r)=J_{0}\,\frac{\pi^{r(d-r)/2}}{\Gamma\Bigl(\frac{r(d-r)}{2}+1\Bigr)}\,c_{2}^{-r(d-r)},$
the claim follows. It remains to show that the Jacobian determinant is bounded below $J_{0}(d,r)>0$. At $A=0$, the chart is an isometry up to second‐order terms, so its differential has singular values all equal to 1. By continuity of the singular values on compact neighborhoods, there exists a small radius $\rho>0$ such that for all $|A|\leq\rho$, the smallest singular value remains at least, $0<\epsilon_{\rho}\leq 1$. Hence, by considering the appropriate value $\rho$ such that $\epsilon_{\rho}=\tfrac{1}{2}$, on $B_{\rho}(0)\subset\mathbb{R}^{r(d-r)}$ the Jacobian determinant is $(\tfrac{1}{2}\bigr)^{r(d-r)}>0$, establishing the desired uniform lower bound. ∎
To ensure that such a subspace $\hat{L}$ preserves the separation between inliers and outliers, we need to understand how perturbations of $L$ affect the distances of points to the subspace. This is precisely the focus of Lemma 3, which bridges the gap between subspace perturbations and the preservation of the inlier-outlier separation. It shows that if $\hat{L}$ is sufficiently close to $L$ then the ordering of distances from points to $\hat{L}$ will preserve the separation between inliers $X\setminus S^{*}$ and outliers $S^{*}$. This result relies on the $\alpha$-gap condition (Lemma 1), which ensures that outliers are sufficiently far from $L$ compared to inliers.
@@
Let $^{\hat{L}}$ be a subspace of rank $r$ in $\mathbb{R}^{d}$ such that
$\|P_{\hat{L}}-P_{L}\|\leq\epsilon\quad\text{with}\quad\epsilon\leq\frac{\alpha}{2}\,\text{dist}_{\min},$
[…]
@@
For any $v\in X$, one has
$|\mathtt{dist}(v,{\hat{L}})-\mathtt{dist}(v,L)|\leq\|P_{\hat{L}}-P_{L}\|.$
Thus, for $p\in X\setminus S^{*}$ and $q\in S^{*}$,
$\mathtt{dist}(p,{\hat{L}})\leq\mathtt{dist}(p,L)+\epsilon,\quad\mathtt{dist}(q,X)\geq\mathtt{dist}(q,L)-\epsilon.$
By the $\alpha$-gap constant (Lemma 1),
$\mathtt{dist}(q,L)-\mathtt{dist}(p,L)\geq\alpha\,\mathtt{dist}(p,L)\geq\alpha\,\text{dist}_{\min}.$
Therefore, if $2\epsilon\leq\alpha\,\text{dist}_{\min},$ then
$\mathtt{dist}(q,L)-\epsilon\geq\mathtt{dist}(p,L)+\epsilon,$
[…]
By combining Lemma 2 and Lemma 3, we can conclude that with a fine probability, a random subspace $\hat{L}$ will preserve the inlier-outlier separation, provided $\epsilon$ is chosen appropriately.
However, Lemma 3 depends critically on the minimum distance $\text{dist}_{\min}$ from the inliers $X\setminus S^{*}$ to the subspace $L$. Specifically, the perturbation threshold $\epsilon=\frac{\alpha}{2}\text{dist}_{\min}$ must be small enough to ensure that the distances of inliers and outliers to $\hat{L}$ are not perturbed too much. This raises the question: what we expect the value of $\text{dist}_{\min}$ to be?
The following lemma answers this question by providing a probabilistic lower bound on $\text{dist}_{\min}$. This result is derived from a geometric argument about the distribution of points in high-dimensional space. The key insight is that, in high dimensions, points are unlikely to cluster too closely around any fixed subspace $L$, and thus $\text{dist}_{\min}$ is bounded below and ensures that the perturbation threshold is sufficiently small to apply Lemma 3, thereby guaranteeing the preservation of the inlier-outlier separation. In particular, we show that with high probability $\mathrm{dist}_{\min}\;\geq\;c_{1}\,n^{-1/(d-r)}$.
@@
Let $S=\{p_{1},\dots,p_{n}\}\subset\mathbb{R}^{d}$ be drawn i.i.d. from the uniform distribution on the unit ball in $\mathbb{R}^{d}$. Let $L\subset\mathbb{R}^{d}$ be any fixed $r$-dimensional subspace, and define
$\mathrm{dist}_{\min}=\min_{p\in S}\mathtt{dist}(p,L)=\min_{p\in S}\bigl\|P_{L^{\perp}}p\bigr\|.$
Then there is a constant $c_{1}=c_{1}(d,r)>0$ such that
$\Pr\bigl[\mathrm{dist}_{\min}\geq c_{1}\,n^{-1/(d-r)}\bigr]\;=\;1-O\bigl(n^{-1}\bigr).$
@@
Since the law of a uniform point $p$ in the $d$-ball is invariant under orthogonal transformations, the marginal distribution of its $L^{\perp}$-component depends only on the radius. Concretely, for any $\delta\in[0,1]$,
$\Pr\bigl(\|P_{L^{\perp}}p\|\leq\delta\bigr)\;=\;\frac{Vol\bigl(B^{d-r}_{\delta}\bigr)}{Vol\bigl(B^{d}_{1}\bigr)}\;=\;\delta^{\,d-r},$
since the volume of a $k$-dimensional Euclidean ball of radius $\delta$ scales like $\delta^{k}$ .
Now we can conclude the result using a Tail bound and union bound. Let $X_{p}=\|P_{L^{\perp}}p\|$. Then $\Pr\bigl(X_{p}\leq\delta\bigr)=\delta^{\,d-r}$. By a union bound over the $n$ independent samples, $\Pr\bigl(\exists\,p\in S:\,X_{p}\leq\delta\bigr)\;\leq\;n\,\delta^{\,d-r}$.
Thus $\Pr\bigl(\mathrm{dist}_{\min}\leq\delta\bigr)\leq n\,\delta^{\,d-r}$. To make this failure probability $\leq n^{-1}$, choose $\delta=c_{1}\,n^{-1/(d-r)}$ with $c_{1}$ a suitable constant, yielding $\Pr(\mathrm{dist}_{\min}\geq\delta)\geq 1-O(n^{-1})$. This completes the proof that, under the stated sampling model, the minimum distance of $S$ to any fixed $r$-subspace $L$ exceeds $c_{1}\,n^{-1/(d-r)}$ with high probability. ∎
It is important to note that the bound holds provided that the number of points $n$ is not large enough to form an adversarial distribution. An adversarial distribution would require at least $n_{\text{max}}=O\left(\delta^{-(d-r)}\right)$ points to densely pack the (dr)(d-r)-dimensional space and break the lower bound on $\text{dist}_{\min}(S,L)$. If $n$ is smaller than this threshold, the points are unlikely to form an adversarial configuration, and the bound on $\text{dist}_{\min}(S,L)$ holds with high probability.
Thus, for $n$ sufficiently small, the minimum distance between the points in $S$ and the subspace $L$ is at least $c_{1}\,n^{-1/(d-r)}$ with high probability.
With the results of Lemmas 2, 3, and 4, in place, we are now ready to derive the probability that a random subspace $\hat{L}$ correctly identifies the inliers $X\setminus S^{*}$ as the (nk)(n-k)-nearest neighbors. This result is significant because it provides a quantitative guarantee for the success of the randomized algorithm 2 in identifying the optimal subspace and removing outliers. The constant $C(d,r,\alpha)$ aggregates the contributions of previously stated lemmas, reflecting the interplay between the geometry of high-dimensional spaces, the separation condition, and the probabilistic nature of the algorithm.
@@
Under assumptions of lemmas 2, 3, and 4, if $^{\hat{L}}$ is chosen uniformly at random from the set of all $r$-dimensional subspace of $\mathbb{R}^{d}$, then
$\Pr\Bigl[X\setminus S^{*}\text{ is the }(n-k)\text{-NN set of }{\hat{L}}\Bigr]\geq C(d,r,\alpha)\,n^{-r}$
where $C$ is a constant depending on $d$, $r$ and $\alpha$. Explicitly, $C(d,r,\alpha)=c_{0}(d,r)\,\left(\frac{\alpha\,c_{1}}{2}\right)^{r(d-r)}.$